OpenStudy (anonymous):

Help! please I so don't understand this problem!?! The area of a play ground is 20 Square yards. The length of the playground is 5 times longer than it's width find the length for the playground length = 1yd, Width = 20 yd length = 2yd, Width = 10 yd length = 10yd, Width = 2 yd length = 20yd, Width = 1 yd

OpenStudy (joemarket):

length = 10yd, Width = 2 yd, because 10 (the length) is 5 times longer than 2 (the width)

OpenStudy (anonymous):

 w=width 5w=length To solve this you have w*5w=20 So solve for w 5w^2=20 Divide by 5 w^2=4 sqrt(w)=sqrt(4) w=2 Then plug in w to find length 5(2)=10 So Length=10 Width=2